# Graph RAG for Legal Norms: A Hierarchical and Temporal Approach

### v2

@@ 3.2 Versions as Aggregations
    groups, directly or indirectly, component C, including the norm (as a whole, as level 0);
@@
In Figure 11, we observe that on the norm’s original date (1988-10-05), the original version of the component "tit2" is formed by the original versions of its children. That is, the text of "Título II” (Title II) on that date consists of it own text and the texts of its children from that same date.
@@
On 1993-09-14, a new version for “tit2 cap4" was produced, either by its own alteration or indirectly by the alteration of one or more components in its tree. A more common implementation would be to produce new versions for the other children of "tit2" on 1993-09-14 and then produce a 1993-09-14 version for "tit2", which would be a composition [^13] In UML terms, Composition represents a strong whole-part relationship where the ”part” belongs exclusively to one ”whole” and shares its lifecycle; if the whole is deleted, the part is too. In our initial comparison, this would mean a parent version strictly owns child versions from the exact same date. of its children’s versions on the same date - similar to the formation of the "tit2" text on the original date. This would imply that, for each date on which the text of the norm was altered, a complete version of the norm would be produced from the composition of the complete versions of each immediate child of its articulation.
@@
However, the other children of "tit2" were not altered on 1993-09-14. Our proposal, therefore, is not to produce versions for the unaltered children (for dates other than the original) and instead reuse the most recent versions of these children. Instead of composition, the 1993-09-14 version for "tit2" would be an aggregation [^14] Aggregation, in UML, represents a weaker whole-part relationship where the ”part” can exist independently of the ”whole” and can potentially be shared by multiple wholes. In our proposed model, this reflects how a child version (the ”part”) can be reused by multiple parent versions (the ”wholes”) across different times without its existence depending on any single one. of the most recent versions of the children.
We can verify the formation of these version aggregations in Figure 11, where new versions of the component "tit2" were produced as a consequence of new versions of one or more of its children on each date, reusing unchanged versions of its children where appropriate.
Figure 11: New versions of the component "tit2" derived from new versions of some of its children
Thus, in the construction of the graph, besides each version being existentially linked to its component, this work proposes that each version, when it is not a leaf, be a agregation with its child versions. In other words, a child version does not "belong" exclusively to a single parent version, but rather contributes to the formation of multiple parent versions at different moments in time.
@@
When a version of the child component Cx does not exist for a date Tn, the most recent version of Cx, prior to Tn, will be used in the aggregation, as exemplified by V1 of C3, which is being used in the formation of version V2 of C1 (examples of this situation also appear in Figure 11). In this case, Version V1 of component C3 contributes to versions V1 and V2 of component C1, demonstrating that a child version can participate in multiple parent versions in different temporal contexts. The orange arrows indicate the aggregation relationship (Figure 12).
@@
Figure 12: Diagram of aggregation relationships between versions, illustrating how child versions can contribute to multiple parent versions at different times and how earlier versions are reused in the absence of subsequent changes.
This representation of versions as aggregation ensures uniqueness in the representation of a version. In a traditional approach, there is a complete version of the articulation tree for each change date. That is, for each component of the tree, there is a version for the date, even if the content of that component was not changed on that date. In the proposed approach, there is a version for the component on the original date and on each date its content was actually changed, thus eliminating redundant and unnecessary copies of the version, contributing to a non-ambiguous and economical representation.
@@
In the context of RAG, it allows the system to retrieve not only the specific version of a component, but also the context of its parent and child versions, providing a more holistic and precise understanding of the norm at a given moment, facilitating more complex queries about the temporal evolution of norms [2].
3.3 Additional Context for Versions
This work also proposes that, for each Version inserted into the graph, information describing its validity status and, when terminated, the type of action that promoted its termination (“revogação” (repeal), “alteração de texto” (text change), “supressão” (suppression), “derrubada de veto” (veto override), among other examples) and the responsible norm (the legal norm, and, ideally, the specific component of that norm, that promoted or commanded the action) be added, when available, as additional context of the version.
This information of a version will be registered by the property Contexto Adicional (Additional Context) of the Version entity in the graph. If new alteration actions occur between subsequent and prior norms, this status can be updated over time.
For example, consider the caput of Art. 6º (presented in Figure 5) of the Brazilian Federal Constitution of 1988, signed and published on 1988-10-05. Thus, at the time of the Constitution’s publication, the Additional Context property of its original version (1988-10-05) will contain information about the start of its validity:
"Essa versão tem vigência de 1988-10-05 até hoje."
("This version is valid from 1988-10-05 until today.")
The version of a component is created using the signature date of its norm (when original) or the signature date of the altering norm, when produced by an action. On the other hand, the validity of the version starts from the publication date (by default in Brazil) [^15] Defining the start of validity for a legal norm and its components is a complex issue, due to a wide variety of possibilities. Therefore, it will not be the focus of this article and will be explored in future work. of the norm (when original) or the publication date of the altering norm, when produced by an action – and the previous version of the component, when exists, has its validity terminated using the date before this publication date.
For example, with the publication, on 2000-02-15, of “Emenda Constitucional nº 26 de 14/02/2000” (Brazilian Constitutional Amendment No. 26), signed on 2000-02-14, an action was commanded, through the caput of its Art. 1º, to alter the wording of the caput provision of Art. 6° of the Brazilian Federal Constitution of 1988 – the action is represented in the graph [^16] The modeling of the “Action presented in Figure 13 was simplified for didactic purposes. It can be improved using ontology techniques, such as those proposed in UFO (Unified Foundational Ontology), found at https://nemo.inf.ufes.br/projetos/ufo/. Accessed on 22.03.2025. by the Action entity (in brown in Figure 13). This action terminated the validity of the 1988-10-05 version and produced a new version, dated 2000-02-14 (with the new text), for the altered component.
For a more precise representation in the Graph, specialized entity types, derived from Action, can be created for each type of action between norms found in the treated legal system. For example, in the Brazilian case, types like "Alteração de Texto" (Text Change Action), "Renumeração" (Renumbering Action), "Revogação" (Repeal Action) etc. Thus, the entity representing the action could be included in the graph using the specialized type.
Figure 13: Representation in the knowledge graph of the text change action
The new version, dated 2000-02-14 (signature date of the Amendment), of the caput of Art. 6º of the Brazilian Federal Constitution of 1988, will then have as the content of the Additional Context, at the time of the Amendment’s publication:
"Essa versão tem vigência de 2000-02-15 até hoje."
("This version is valid from 2000-02-15 until today.")
Furthermore, the content of the Additional Context of the previous version (dated 1988-10-05) of the caput of Art. 6° of the Brazilian Federal Constitution of 1988 will be updated, with a complement textually describing the reason for and the entity responsible for the termination of its validity:
"Essa versão tem vigência de 1988-10-05 a 2000-02-14. Sua vigência foi encerrada pela nova redação de texto dada a esse dispositivo pelo caput do Art. 1º da Emenda Constitucional nº 26 de 14/02/2000."
("This version was valid from 1988-10-05 to 2000-02-14. Its validity was terminated by the new text wording given to this component by the caput of Art. 1º of Constitutional Amendment No. 26")
The additional contexts can be persisted directly as textual metadata (property) of the version or be generated automatically from the actions (and from start-of-validity and end-of-validity information) inserted in the graph.
Either way, the content of the Additional Context should not be used in the creation of the Text Unit embedding of its version, so as not to add noise to its semantic understanding.
The Additional Context is important to enrich the response generated by the LLM in the RAG process. When the RAG system retrieves a specific Version as relevant context, the Additional Context associated with this version can be sent along with the Text Unit to the LLM, providing valuable information about the history and validity of the component, allowing for more complete and contextually rich responses – in more advanced implementations, the additional context would be sent to the LLM only when it is sensed that it is necessary for generating the response. This will provide context for the LLM to display, for example, besides the texts in the timeline of a component, also information about the validity of each text and, in case of termination, by whom and for what reason its validity was terminated.
3.4 Alteration Actions as Text Units
As previously presented, a subsequent norm can mandate an amendment to the wording of a provision of a prior norm, a repeal of a provision, or of a prior norm in its entirely, etc. An additional proposal, therefore, is to create a Text Unit representing each alteration action between legal norms/components. Considering the action illustrated in Figure 13, the content of the new Text Unit will be:
@@ 3.4 Alteration Actions as Text Units
"A Emenda Constitucional nº 26, de 14 de fevereiro de 2000, por meio do caput do seu Art. 1º, deu uma nova redação ao caput do Art. 6º da Constituição Federal de 1988. Essa alteração encerrou em 2000-02-14 a vigência da versão original desse dispositivo (de 1988-10-05) e instituiu uma nova versão com vigência a partir de 2000-02-15, cujo texto passou a ser: ’São direitos sociais a educação, a saúde, o trabalho, a moradia, o lazer, a segurança, a previdência social, a proteção à maternidade e à infância, a assistência aos desamparados, na forma desta Constituição.”
@@
Although this alteration command already appears in the text of the caput of Art. 1º of Constitutional Amendment no. 26 (amending component/norm), as shown in Figure 14, generating a Text Unit for the Action entity structures and completes the information about the action/command, making explicit, for example, the validity periods of the text versions.
Figure 14: Text of the Emenda Constitucional nº 26 de 14/02/2000 (Constitutional Amendment no. 26)
The generated Text Unit is then associated with the Action entity in the graph. And it receives its own embedding, allowing, besides Version Text Units embeddings, for more sophisticated analyses of legislative evolution, facilitates the traceability of normative alterations, and improves the understanding of the life cycle of legal norms in the knowledge graph in the context of RAG [14].
@@ 3.5 Metadata as Text Units
Besides the texts of legal norms, other sources of information can be used to feed the knowledge graph. One possibility would be to generate Text Units with the metadata [^17] Metadata is data about data. In this context, structured or textual information describing the properties and characteristics of legal norms, such as publication date, authorship, type of norm, relationships with organizations, etc., in textual format, to associate them with the Norm and Component entities, similar to the original proposal of entity summaries from Graph RAG – the difference here is that this information, generally, is not contained within the text of the norm itself; therefore, these summaries could not be extracted from the text chunks, as in the original proposal.
@@
As example, for the Norm entity representing the legal norm “Lei nº 10.406 de 10/01/2002” (Law no. 10,406 of 2002-01-10), a Text Unit would be generated textually describing the entity’s properties [^18] This Text Unit was generated from the structured metadata registered in the knowledge graph of the ”Sistema de Gestão de Normas Jurídicas (Sigen)” (Legal Norm Management System) of the Federal Senate of Brazil. This metadata can also be viewed publicly, in structured form, through the Normas.leg.br portal, as in the example: https://normas.leg.br/?urn=urn:lex:br:federal:lei:2002-01-10;10406. Accessed on 2025-03-22.:
@@
"São apresentados a seguir os metadados gerais da norma jurídica federal ’Lei nº 10.406 de 10/01/2002’: é do tipo ’Lei Ordinária’; está no nível ’Legal’ do ordenamento jurídico brasileiro; possui como urn: ’urn:lex:br:federal:lei:2002-01-10;10406’; possui como URL principal (uri): ’https://normas.leg.br/?urn=urn:lex:br:federal:lei:2002-01-10;10406’; possui como URLs alternativas: ’http://www.lexml.gov.br/urn/urn:lex:br:federal:lei:2002-01-10;10406’ e ’http://legis.senado.leg.br/legislacao/DetalhaSigen.action?id=552282’; foi apreciada e aprovada pelo ’Congresso Nacional’; foi assinada em Brasília, 10 de janeiro de 2002; 181º da Independência e 114º da República; foi assinada por FERNANDO HENRIQUE CARDOSO, Aloysio Nunes Ferreira Filho; foi oficialmente publicada pela ’Imprensa Nacional’, tendo como data de publicação original e oficial: 11/01/2002; tem como nomes alternativos: ’LEI-10406-2002-01-10’ e ’Código Civil (2002) (CC)’; tem como ementa: ’Institui o Código Civil’."
@@
("The general metadata for the federal legal norm Law no. 10,406 of 2002-01-10 are presented below: it is of the type ’Ordinary Law’; it is at the ’Legal’ level of the Brazilian legal system; its URN is: ’urn:lex:br:federal:lei:2002-01-10;10406’; its main URL (URI) is: ’https://normas.leg.br/?urn=urn:lex:br:federal:lei:2002-01-10;10406’; its alternative URLs are: ’http://www.lexml.gov.br/urn/urn:lex:br:federal:lei:2002-01-10;10406’ and ’http://legis.senado.leg.br/legislacao/DetalhaSigen.action?id=552282’; it was reviewed and approved by the ’National Congress’; it was signed in Brasília, on January 10, 2002; 181st year of Independence and 114th year of the Republic; it was signed by FERNANDO HENRIQUE CARDOSO, Aloysio Nunes Ferreira Filho; it was officially published by the ’National Press’, with the original and official publication date being: 2002-01-11; it has alternative names: ’LEI-10406-2002-01-10’ and ’Civil Code (2002) (CC)’; its summary is: ’Institutes the Civil Code’.")
@@
Informative relationships between legal norms, such as regulation, succession, correlation, etc., do not promote actions that alter the text of a norm, thus not generating temporal versions. Nevertheless, they are important in the context of semantic retrieval of legal information. Therefore, we could also generate Text Units with the textual descriptions of these relationships.
@@
Considering the legal norm ’CONSTITUIÇÃO DO BRASIL (1967)’ (Constitution of Brazil (1967)), the following are the Text Units [^19] It is not necessary to generate a Text Unit for the other side of the relation, since an LLM and a RAG retrieval already semantically understand this reversal (”succeeded” vs “was succeeded by”, for example). for its relationships presented in the graph of Figure 15.
@@
Figure 15: Example representation of inter-normative relationships (succession) between different normative acts in the knowledge graph.
@@
This approach aligns with the concept of “Multi-aspect embeddings[19], which proposes using multiple vectors to represent information at different levels of detail or different aspects. Instead of generating a single, complete Text Unit for an entity encompassing all its properties and relationships, this work proposes generating one Text Unit for the entity’s intrinsic properties (general metadata), as exemplified above, and a separate Text Unit for each of its relationships.
Thus, instead of having just one vector/point in the vector space representing an entity, we will have multiple vectors/points associated with it. Each vector will capture/encode a different aspect of the information related to the entity (potentially viewed as a different "lens" through which the entity is represented), thus offering a more contextualized and granular view of the entity, which can lead to significant improvements in semantic retrieval by proximity.
It is also proposed to generate Text Units for Component entities. For example:
@@
These metadata Text Units about the norm, or component, can be generated externally (from other information sources) and inserted directly into the graph, or be generated from the properties and/or relationships of the entity in the graph [^20] As these properties (publication date, alternative names, level in the legal system, etc.) and relationships (correlation, succession, etc.) generally cannot be directly derived from the texts of the legal norms, they must be populated in the graph either manually or through semi-automatic extraction processes using other information sources. - although the original metadata might be structured in the graph, converting them to textual format allows their uniform integration and processing within the Graph RAG workflow.
The metadata Text Unit for a norm or component will have its own embedding generated and will also be subject to semantic retrieval, like the content Text Units derived from versions and alteration actions. Consequently, vector search will be able to retrieve information about the metadata (properties and relationships) of a legal norm (or a component), even though this information is generally not present in the text of the norm itself.
We can see these Text Units derived from metadata, additional to those derived from versions of the norms/components, in Figure 16. The insertion of this type of Text Unit brings additional information to the Knowledge Graph and, consequently, to the vector database of contexts.
Figure 16: Knowledge graph illustrating the Text Units derived from both versions of the norms/components and their metadata.

### v3

@@ 3.2 Temporal Versions as Aggregations
    groups, directly or indirectly, component C, including the norm as a whole (as level 0; in this case, it is a TV);
@@
This propagation of change is where the concept of aggregation becomes critical. Consider Figure 11. On the norm’s original date (1988-10-05), the original CTV of the component "tit2" is a composition of the original CTVs of all its children from that same date.
@@
On 1993-09-14, a new CTV for component “tit2 cap4” was created due to an amendment. A naive approach would be to create new 1993-09-14 CTVs for all other children of "tit2" (even the unchanged ones) and then compose them to form a new CTV for "tit2". This would be a Composition [^13] In UML terms, Composition represents a strong whole-part relationship where the ”part” belongs exclusively to one ”whole” and shares its lifecycle; if the whole is deleted, the part is too. In our initial comparison, this would mean a parent version strictly owns child versions from the exact same date.. This approach is inefficient, creating redundant data and obscuring which components actually changed.
@@
Our proposal, however, is to model the new 1993-09-14 CTV for "tit2" as an Aggregation [^14] Aggregation, in UML, represents a weaker whole-part relationship where the ”part” can exist independently of the ”whole” and can potentially be shared by multiple wholes. In our proposed model, this reflects how a child version (the ”part”) can be reused by multiple parent versions (the ”wholes”) across different times without its existence depending on any single one. of its children’s most recent CTVs. This means the new CTV for "tit2" is formed by the new CTV of "tit2 cap4" and the reused, pre-existing CTVs of all its other, unaltered children.
This aggregation logic is shown in Figure 11. New CTVs for "tit2" are created only when one or more of its children are amended, and these new parent CTVs simply point to the latest available CTV of each child.
Figure 11: New Temporal Versions of the component "tit2" derived from new CTVs of some of its children. Unchanged child components have their most recent CTV reused (e.g., the 1988-10-05 CTV of tit2_cap2 is aggregated into the 1993-09-14 CTV of tit2).
@@
This model establishes that a child’s Temporal Version does not "belong" exclusively to a single parent CTV but can be aggregated into multiple parent CTVs across different points in time, as illustrated in Figure 12. When retrieving the state of a component Cx for a date Dn for which no CTV exists, the system simply traverses back in time to find the most recent CTV prior to Dn. It is exemplified by the Version V1 of component C3, which contributes to the versions V1 and V2 of component C1, demonstrating that a child version can participate in multiple parent versions in different temporal contexts. The orange arrows indicate the aggregation relationship.
@@
Figure 12: Diagram of aggregation relationships between Temporal Versions, illustrating how child CTVs can be reused by multiple parent CTVs at different times.
This representation of Temporal Versions as aggregations ensures uniqueness and efficiency. Instead of creating a complete version of the entire document tree for each amendment date, we only create new CTV nodes for components whose text (or whose descendants’ text) has actually changed. This eliminates redundancy, provides a non-ambiguous and economical representation, and accurately reflects the real-world legislative process.
@@
In the context of RAG, this allows the system to retrieve not only the specific CTV (and its CLV/Text Unit) of a component, but also the precise context of its parent and child CTVs for any given moment, facilitating complex queries about the temporal evolution of norms [2].
3.3 Modeling Normative Events as Actions
The lifecycle of a Temporal Version (TV/CTV) — its creation and termination — is governed by normative events, such as amendments, repeals, or original enactments. To capture this causality explicitly, we introduce an Action entity into our graph. This approach is grounded in event-centric ontologies for the legal domain, which model legislative changes as discrete Events that occur at a specific time and trigger consequences [18].
An Action node represents a specific command within a legal norm that affects another norm or component. Examples include text amendments, repeals, suppressions, or renumberings. As illustrated in Figure 13, an Action has a dual role:
  1. 1.
    It terminates the validity of a prior Temporal Version.
  2. 2.
    It produces a new Temporal Version.
It can link the CTV (and the component) of the amending norm (the cause) to both the old TV/CTV that it terminates and the new TV/CTV that it creates.
Figure 13: Representation of a legislative Action (Event) in the knowledge graph. The Action, commanded by the Art.1’s caput of the Brazilian Constitutional Amendment 26, terminates the validity of the original 1988-10-5 CTV of Art. 6’s caput of the Brazilian Federal Constitution of 1988 and produces its new 2000-02-14 CTV.
For a more precise representation, specialized entity types derived from Action can be created, such as "Text Change Action" or "Repeal Action".
To make these events semantically retrievable, we propose creating a descriptive Text Unit for each Action. This Text Unit is a natural language summary of the action, its cause, and its effects. For the action illustrated in Figure 13, the content of this new Text Unit would be:
@@ 3.3 Modeling Normative Events as Actions
"A Emenda Constitucional nº 26, de 14 de fevereiro de 2000, por meio do caput do seu Art. 1º, deu uma nova redação ao caput do Art. 6º da Constituição Federal de 1988. Essa alteração encerrou em 2000-02-14 a vigência da versão original desse dispositivo (de 1988-10-05) e instituiu uma nova versão com vigência a partir de 2000-02-15, cujo texto passou a ser: ’São direitos sociais a educação, a saúde, o trabalho, a moradia, o lazer, a segurança, a previdência social, a proteção à maternidade e à infância, a assistência aos desamparados, na forma desta Constituição.”  [^15] The validity began on 2000-02-15 because the Amendment was published on that date, and its text stipulated: ’This Constitutional Amendment enters into force on the date of its publication.
@@
Although this alteration command is implicitly present in the text of the amending norm (Figure 14), generating a dedicated Text Unit for the Action entity makes the event’s details explicit and directly queryable. It structures and completes the information about the command, clarifying, for example, the validity periods of the affected TV/CTVs.
Figure 14: Text of the Emenda Constitucional nº 26 de 14/02/2000 (Constitutional Amendment no. 26), which contains the command for the alteration action.
This generated Text Unit is then associated with the Action entity in the graph. Since this content is a linguistic description, it can exist in multiple languages, with each translation being a distinct Text Unit of the same Action node.
Like other Text Units, it receives its own embedding. This allows for sophisticated semantic searches not only on the content of legal norms (CLVs Text Units) but also on the events that shape their evolution (Actions Text Units), facilitating the traceability of normative alterations and improving the understanding of the life cycle of legal norms in the context of RAG [19].
Furthermore, these Action Text Units are also crucial for enriching the response generated by the LLM in the RAG process. When the system retrieves a specific Language Version as relevant context, the Text Unit of the Action that created or terminated its corresponding Temporal Version can also be provided to the LLM. This gives the model valuable information about the history and validity of the component, allowing for more complete and contextually rich responses. For instance, the LLM could then generate an answer that not only provides the text of a provision but also explains when it came into force, when it was repealed, and by which specific legislative act it was altered.
@@ 3.4 Metadata as Text Units
Besides the textual content of legal norms captured in Language Versions, other sources of information can be used to feed the knowledge graph. A proposed technique is to generate Text Units from structured metadata [^16] Metadata is data about data. In this context, structured or textual information describing the properties and characteristics of legal norms, such as publication date, authorship, type of norm, relationships with organizations, etc.. These Text Units are textual summaries associated directly with the abstract Norm and Component entities (the Works). This is similar to the entity summaries in the original Graph RAG proposal, but with a key difference: this information is generally not contained within the norm’s text itself and must be sourced externally or from graph properties.
@@
For example, for the Norm entity representing the legal work “Lei nº 10.406 de 10/01/2002”, a metadata Text Unit would be generated to textually describe its properties [^17] This Text Unit was generated from the structured metadata registered in the knowledge graph of the ”Sistema de Gestão de Normas Jurídicas (Sigen)” (Legal Norm Management System) of the Federal Senate of Brazil. This metadata can also be viewed publicly, in structured form, through the Normas.leg.br portal, as in the example: https://normas.leg.br/?urn=urn:lex:br:federal:lei:2002-01-10;10406. Accessed on 2025-03-22. - like any textual description, this Text Unit can be generated in multiple languages.
@@
"São apresentados a seguir os metadados gerais da norma jurídica federal ’Lei nº 10.406 de 10/01/2002’: é do tipo ’Lei Ordinária’; está no nível ’Legal’ do ordenamento jurídico brasileiro; possui como urn: ’urn:lex:br:federal:lei:2002-01-10;10406’; possui como URL principal (uri): ’https://normas.leg.br/?urn=urn:lex:br:federal:lei:2002-01-10;10406’; possui como URL alternativa: ’http://www.lexml.gov.br/urn/urn:lex:br:federal:lei:2002-01-10;10406’; foi apreciada e aprovada pelo ’Congresso Nacional’; foi assinada em Brasília, 10 de janeiro de 2002; 181º da Independência e 114º da República; foi assinada por FERNANDO HENRIQUE CARDOSO, Aloysio Nunes Ferreira Filho; foi oficialmente publicada pela ’Imprensa Nacional’, tendo como data de publicação original e oficial: 11/01/2002; tem como nomes alternativos: ’LEI-10406-2002-01-10’ e ’Código Civil (2002) (CC)’; tem como ementa: ’Institui o Código Civil’."
@@
("The general metadata for the federal legal norm Law no. 10,406 of 2002-01-10 are presented below: it is of the type ’Ordinary Law’; it is at the ’Legal’ level of the Brazilian legal system; its URN is: ’urn:lex:br:federal:lei:2002-01-10;10406’; its main URL (URI) is: ’https://normas.leg.br/?urn=urn:lex:br:federal:lei:2002-01-10;10406’; its alternative URL is: ’http://www.lexml.gov.br/urn/urn:lex:br:federal:lei:2002-01-10;10406’; it was reviewed and approved by the ’National Congress’; it was signed in Brasília, on January 10, 2002; 181st year of Independence and 114th year of the Republic; it was signed by FERNANDO HENRIQUE CARDOSO, Aloysio Nunes Ferreira Filho; it was officially published by the ’National Press’, with the original and official publication date being: 2002-01-11; it has alternative names: ’LEI-10406-2002-01-10’ and ’Civil Code (2002) (CC)’; its summary is: ’Institutes the Civil Code’.")
@@
Informative relationships between legal norms, such as regulation, succession, or correlation, are also modeled as relationships between Norm entities (Works). These relationships do not generate new temporal versions but are fundamental for semantic retrieval. We can therefore generate Text Units to describe them.
@@
Considering the legal norm ’Constituição do Brasil (1967)’ (Brazilian Constitution of 1967), the following are the Text Units [^18] It is not necessary to generate a Text Unit for the other side of the relation, since an LLM and a RAG retrieval already semantically understand this reversal (”succeeded” vs “was succeeded by”, for example). for its succession relationships, as shown in Figure 15.
@@
Figure 15: Example representation of inter-normative relationships (succession) between different Norm (Work) entities in the knowledge graph.
@@
Text Units can also be created for relationships involving Component entities, for example:
@@
This approach aligns with the concept of “Multi-aspect embeddings[20], which proposes using multiple vectors to represent different aspects of an entity. A Norm or Component (Work) can have multiple associated Text Units (including, in various languages):
  • One Text Unit for its intrinsic properties (metadata summary).
  • A separate Text Unit for each of its relationships (e.g., succession, correlation).
Thus, instead of a single vector, an abstract entity like a Norm will be associated with multiple vectors/points in the vector space. Each vector captures a different "lens" on the entity, enabling more contextualized and granular semantic retrieval.
These metadata Text Units can be generated externally or from the structured properties and relationships already present in the graph [^19] As these properties (publication date, alternative names, level in the legal system, etc.) and relationships (correlation, succession, etc.) generally cannot be directly derived from the texts of the legal norms, they must be populated in the graph either manually or through semi-automatic extraction processes using other information sources.. Converting them to a textual format allows for their uniform processing within the Graph RAG workflow.
Each metadata Text Unit will have its own embedding. Consequently, vector search can retrieve information about a Norm or Component based on its metadata, even if that information is not present in its textual content. Figure 16 illustrates this concept, showing how Text Units can be derived both from Language Versions (content of the compiled legal norm) and from the abstract Norm/Component (and Action) entities themselves (metadata).
Figure 16: Knowledge graph illustrating how Text Units are derived from two sources: from Language Versions (representing content) and from the abstract Norm/Component (and Action) entities (representing metadata).